OpenStudy (anonymous):

A carpenter is building a rectangle room with a fixed perimeter of 172ft. What dimensions would yield the maximum area? What is the maximum area? The length that would yield the maximum area is ? ft

OpenStudy (anonymous):

perimeter of a rectangle is to add up the lengths of the sides so since there are 2 lengths n 2 widths then u kno that if we called the width w and the length l then the perimeter equals 2 times the addition of the length and the width: 2(l+w)=perimeter and u kno the peremeter is 172 so: 2(l+w)=172 Now the are equals length times width so: A=lw now solve the first equation for l 2(l+w)=172 l+w=86 l=86-w substitute this l for the l in the 2nd equation A=lw A=(86-w)w A=86w-w^2 Now find the maximum w=-(86)/(2*-1) w=43 so l=43 and the best dimensions is a square with length 43

OpenStudy (anonymous):

43 feet is the answer

OpenStudy (anonymous):

then i did it right :)

OpenStudy (anonymous):

wide?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what is the width? I tried 86 but that was not right

OpenStudy (anonymous):

its a square. the length and the width r the same. l=w=43

OpenStudy (anonymous):

maximum area? please

OpenStudy (anonymous):

43*43 sq ft.